OpenStudy (anonymous):

(FoG)(x) f(x)=x^2-3+2 g(x)=x-2

OpenStudy (anonymous):

in f(x) in place of every x put x-2 ..

OpenStudy (anonymous):

i got x^3-3x^2+2x-8 is that correct

OpenStudy (anonymous):

x^3-2x^2-3x-4

OpenStudy (anonymous):

no its .. i think the ques is f(x)= x^2 -3x +2 then fog = x^2 -7x +12

OpenStudy (anonymous):

cool I got it. I was substituting incorrectly

OpenStudy (anonymous):

hmm gud..